OpenStudy (xapproachesinfinity):

How is Coulomb Law verified How did Coulomb arrived to the equation \[F=K\frac{Q_1 Q_2}{r^2}\]

candycove (candycove):

:D

OpenStudy (xapproachesinfinity):

hehe ;)

candycove (candycove):

:P

imqwerty (imqwerty):

verified through experiments nd its not always valid its valid for only static and point charges and \(F ∝ Q_1Q_2\) \(F ∝\large \frac{1}{r^2}\) so \(F=K\large \frac{Q_1 Q_2}{r^2}\)

OpenStudy (xapproachesinfinity):

how it is checked on point charges! to begin with what is meant by a point charge how do you separate a single charge from a materiel and experiment on it

imqwerty (imqwerty):

we can prove that its not valid for charges which are not point charges for example take this case-  now when + charge moves to left and charge density increases on left and in \(F=k\frac{Q_1Q_2}{r^2}\) the r that must be taken will be greater than our r as the overall charge moved left so it wo't work moreover we can say that when a electric force is applied then the body dipoles or distribution of charge may change due to polarization and it will affect the net force acting on body